OpenStudy (anonymous):

does anyone understand this stuff? A local Radio Shack Store wants to buy a new line of Plasma TV's. Manufacturer A offers chain discounts of 18/12, while Manufacturer B offers terms of 17/13. The single equivalent discount rate of ? is the best deal.

OpenStudy (anonymous):

so confused

OpenStudy (anonymous):

A dicount offered = (100-18)(100-12)/(100*100) B discount offered = (100-17)(100-13)/(100*100) calculate to get answer

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

aah you are doing it wrong

OpenStudy (anonymous):

total discount offered by any manufecturer =(100-D1)/100 * (100-D2)/100 and so on for example if firstly someone offer 30% discount then 40% that means (100-30)/100* (100-40)/100=0.70 * 0.60 =0.42 so total discount offered is " (100-42)=58% " rather than 30%+40%=70% now solve it for your two cases and then compare them. . .

OpenStudy (anonymous):

A LOCAL RADIO SHACK WANTS TO BUY A NEW LINE OF PLASMA TT,UF THE MANUFACTURE OFFERS A CAHIN DISCOUNT OF 18/12? wHATIS THE SINGLE EQUIVlent rate